I need help with money and compound interest for example principal $8500 rate 7% compounding quartley for 5 years I am not sure how to set up and solve sets, also I am gettin confused with investments like finding the accumulated value of an investment for instance investment of $2000 at 8% compounded semi annually for 9 years
Answer by rfer(12670)

A=P(1+r/4)^nt
A=8500(1+0.07/4)^20
A=8500(1.0175)^20
A=8500(1.4148)
A=$12,025.80
-----------------
same thing only divide rate by 2.
A=2000(1+0.08/2)^18
A=2000(1.04)^18
A=2000(2.0258)
A=$4,051.60